OpenStudy (anonymous):

can someone please show me how to find the integral of y^2/25+y^2

OpenStudy (anonymous):

do you mean y^2/(25+y^2)? or (y^2/25) + y^2?

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

ok, did you get past the long division at the beginning which gives you 1- (25/y^2+25)?

OpenStudy (anonymous):

why do i have to use long division

OpenStudy (anonymous):

i would, you have same degree of y in numerator & denominator.

OpenStudy (anonymous):

oh yes you are correct

OpenStudy (anonymous):

long division gives 1 - (25/(y^2+25))...integral of 1 = y, integral of -25/(y^2+25) = -5arctan(y/5).....so the solution is y-5arctan(y/5)...

OpenStudy (anonymous):

can you please explain how you got the integral of -25/(y^2+25)

OpenStudy (anonymous):

25 * (1/(Y^2 + 5^2))--once you see the trig integrals enough, they just become easy to recognize.

OpenStudy (anonymous):

integral of 1/(x^2 + a^2) = (1/a)arctan(x/a)

OpenStudy (anonymous):

okay thank you